OpenStudy (anonymous):

Find the missing side length of each right triangle. Round to the nearest tenth. 1. a = 12, b = 9, c = ? 2. a = 14, b = ?, c = 35 3. a = ?, b = 4, c = 41.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (angela210793):

u can use Pythagoras Theorem

OpenStudy (anonymous):

I don't really get that. Can you help me with these 3 problems?

OpenStudy (anonymous):

a^2 + b^2 = c^2 is your basic equation to work with. Each problem gives you two of the three variables. Substitute what's given into the equation and solve for the remaining variable.

OpenStudy (angela210793):

1)c^2=a^2-b^2=144-81=63 c=sqr63=close to 8 2)b^2=c^2-a^2=1225-196=1029 b=sqr1029=32 3)a^2=c^2-b^2=1681-16=1665 a=sqr1665=40 also check this link...it may help http://en.wikipedia.org/wiki/Pythagorean_theorem

OpenStudy (anonymous):

144 - 81 =63 which is nearly equal to 8 so the c may be 8 1225 - 196 = 1029 which is equal to 32 so the b is 32 1681 - 16 =1665 which is equal to 40 so the a is 40 if u like this answer plz click the good answer button